# Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs

### v1

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have enhanced their versatility as they integrate a growing number of modalities. Considering the heavy cost of training MLLMs, it is necessary to reuse the existing ones and further extend them to more modalities through Modality-incremental Continual Learning (MCL). However, this often comes with a performance degradation in the previously learned modalities. In this work, we revisit the MCL and investigate a more severe issue it faces in contrast to traditional continual learning, that its degradation comes not only from catastrophic forgetting but also from the misalignment between the modality-agnostic and modality-specific components. To address this problem, we propose an elegantly simple MCL paradigm called "MErge then ReAlign" (MERA). Our method avoids introducing heavy training overhead or modifying the model architecture, hence is easy to deploy and highly reusable in the MLLM community. Extensive experiments demonstrate that, despite the simplicity of MERA, it shows impressive performance, holding up to a 99.84% Backward Relative Gain when extending to four modalities, achieving a nearly lossless MCL performance.
@@
Dingkun Zhang1, Shuhan Qi1thanks: Corresponding author., Xinyu Xiao1, Kehai Chen1, Xuan Wang1, 1Harbin Institute of Technology, Shenzhen dingkunzhang0xffff@gmail.com, {shuhanqi,wangxuan}@cs.hitsz.edu.cn, 23b951021@stu.hit.edu.cn, chenkehai@hit.edu.cn
@@ 1 Introduction
With the recent trend of developing general-purpose any-modality Multimodal Large Language Models (MLLMs)(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024), MLLMs are evolving towards integrating more modalities. The typical MLLM architecture includes modality-specific encoders, modality-specific connectors, and a shared Large Language Model (LLM). A standard process of training MLLMs involves aligning modality-specific components with LLM through modality-text paired data and then fine-tuning on modality-text instruction data. Such architecture and training strategy have been successfully applied to a wide range of modalities, i.e., image(Liu et al., 2024b, a), video(Lin et al., 2024a; Maaz et al., 2024), audio(Li et al., 2024; Wu et al., 2024), point cloud(Chen et al., 2024), etc, equipping MLLMs with the ability to understand a growing number of modalities.
Existing methods(Wu et al., 2024; Zhan et al., 2024; Panagopoulou et al., 2023; Fu et al., 2024) typically employ a joint training strategy, where the MLLM is simultaneously trained on datasets of all predefined modalities. However, it is challenging to extend an existing MLLM to new modalities.
Accordingly, Continual Learning (CL) is proposed to learn from a stream of data. When incrementally learning on the new data, a phenomenon called catastrophic forgetting(McCloskey and Cohen, 1989; Goodfellow et al., 2014) often occurs, i.e., the model forgets the previously learned knowledge. To this end, many CL methods(Kirkpatrick et al., 2017; Yu et al., 2024b; Scialom et al., 2022; Wang et al., 2024) have been proposed to alleviate catastrophic forgetting. Moreover, Modality-incremental Continual Learning (MCL)(Yu et al., 2024a) focuses on the particular scenario of incrementally extending MLLMs to new modalities. However, it is worth noting that: the performance degradation encountered in MCL comes not only from forgetting but also from the misalignment between modality-agnostic and modality-specific components.
@@
The first stage of MERA aims at addressing the forgetting problem. Inspired by the great success of model merging technique in multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a), we introduce model merging to our MCL framework for the purpose of mitigating forgetting. In this work, we focus on the simplest model merging method, i.e., weight averaging, and revise it into an MCL form, aiming to provide a basic framework. We achieve this by associating its merging coefficients with the progress of CL stages and only merging the modality-agnostic components.
The second stage of MERA aims at addressing the misalignment problem. We leverage a small subset of data from each learned modality to realign the modality encoders with the LLM backbone. In this stage, modality encoders and LLM backbone are both frozen, only the connectors are updated to enable an efficient realignment between them. Further experiments show that the realigning stage can significantly narrow the gap between the incrementally learned MLLM and the individually trained expert MLLMs on each modality.
@@
  • We revisit the Modality-incremental Continual Learning (MCL) and investigate a more severe issue it faces in contrast to traditional continual learning: its performance degradation comes not only from forgetting but also from misalignment.
@@
  • Extensive experiments demonstrate that our MERA significantly outperforms other representative continual learning methods, and even achieves a nearly lossless MCL performance.
@@
Table 1: Characteristics of different CL categories and our proposed MERA. Extra-train-memory-free: whether introduces extra GPU memory overhead at training time. Arch-modification-free: whether requires modifying the architecture of the model or adding auxiliary components. Replay-data-free: whether requires access to partial data from the previous tasks or distributions. $\bullet$ denotes that some methods of this category don’t satisfy the property. denotes that this drawback is relatively minor in real applications.
@@ 2.1 Multimodal Large Language Models
Recent advances(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024) in Multimodal Large Language Models (MLLMs) have extended Large Language Models (LLMs) to perceive multimodal inputs such as image, video, audio, point cloud, etc. Early attempts like Flamingo(Alayrac et al., 2022) integrate vision encoders with LLMs via cross-attention to perform image captioning and visual question answering (VQA) tasks. Subsequent methods like InstructBLIP(Dai et al., 2023) leverage instruction tuning to build a general-purpose multimodal system. These attempts lead to the rapid development of MLLMs.
@@
Among these MLLMs, the most influential one is LLaVA(Liu et al., 2024b, a), which utilizes a simple MLP connector to project visual information encoded by the pre-trained vision encoder into the language embedding space while leveraging visual instruction tuning. Due to its simplicity and effectiveness, LLaVA-like architecture is widely adopted by a wide range of subsequent MLLMs(Lin et al., 2024b, a; Maaz et al., 2024; Wu et al., 2024; Chen et al., 2024). In this paper, we assume that the MLLM has a LLaVA-like architecture that includes modality-specific encoders and connectors, and a shared modality-agnostic LLM.
2.2 Continual Learning
@@
Extending existing MLLMs to new modalities can be viewed as a Continual Learning (CL) problem. CL approaches allow models to continually acquire new knowledge with minor forgetting of previously learned knowledge. Existing CL methods mainly fall into the following three categories: Regularization-based methods(Kirkpatrick et al., 2017; Huszár, 2017; Schwarz et al., 2018) seek to protect the parameters that store important knowledge. However, storing the importance matrix requires extra memory with the same scale of the trainable parameters at training time. Architecture-based methods(Yu et al., 2024a, b; Zadouri et al., 2024) add task-specific parameters to the base model for each new task. The drawbacks of this category are that it requires modifications of the model architecture, harming its reusability, and the model scale often grows linearly as tasks increase, introducing extra memory overhead. Replay-based methods(Scialom et al., 2022; Wang et al., 2024) leverage a small subset of previous data and replay them while learning on new data. A drawback of this category is that it requires access to partial data from the previous tasks or distributions. However, this drawback is relatively minor in real applications as the replay data are often accessible.
Table 1 summarizes the characteristics of different CL categories. Among all the CL methods, replay-based methods are the simplest and most widely used for LLMs since they are extra-train-memory-free and arch-modification-free. In this work, we propose a simple MCL paradigm that has the same advantages as replay-based methods.
2.3 Model Merging
Model merging is an emerging technique often used for multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a). It aims to combine the strengths of multiple isomorphic models into one unified model. The simplest model merging method is weight averaging(Utans, 1996), where weights from each model are averaged to form the merged model. Model merging is based on the assumption that the models being merged lie in the same flat loss basin(Neyshabur et al., 2020), hence their interpolations also have low losses. Aside from its applications in multi-task learning, model merging can be naturally leveraged to reduce forgetting(Marczak et al., 2024).
3 Analysis of Degradations in MCL
@@
Modality-incremental Continual Learning (MCL) is a special scenario of CL, where models incrementally learn on the data from new modalities. The first attempt of MCL is (Yu et al., 2024a). However, we point out that the MCL faces a more severe problem. In contrast to traditional CL scenarios, where the degradation comes solely from the forgetting of old knowledge, the degradation encountered in MCL comes from two aspects: forgetting and misalignment.
Forgetting: the modality-agnostic components (the LLM backbone for MLLM) forget the knowledge of old modalities.
Misalignment: the modality-agnostic components are misaligned with the modality-specific ones. When adapting to a new modality, the LLM backbone and the newly added modality encoder and connector are updated while the old ones are kept frozen. Therefore, the LLM backbone inevitably drifts away from the original multimodal feature space of old modality encoders (even the feature space of the new modality encoder, if its connector is not synchronously updated with LLM backbone). This process results in misalignment, further worsening the degradation. A potential exception to misalignment is when replay-based methods are applied since the old modality-specific components are also updated synchronously. However, empirical results in Section 5.5 suggest that replay-based methods still suffer from a certain degree of misalignment and it can be compensated by our proposed realigning stage.
@@
Due to the existence of misalignment, MCL problem requires special treatments compared to traditional CL problems.
@@
4 Method
First, we define the Modality-incremental Continual Learning (MCL) problem as follows. Given a sequence of $m$ modalities $\{M_{1},M_{2},\dots,M_{m}\}$ and their corresponding datasets $\{D_{1},D_{2},\dots,D_{m}\}$, the model sequentially learns on each $D_{i}$ with the permission to access a $r\%$ size subset of previous data as replay data [^1] Different from the definition of replay data in the context of replay-based methods, where $R_{i}$ is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$. In this paper, if not explained, $R_{i}$ for replay-based methods is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$ while $R_{i}$ for our realigning stage is sampled from $\{D_{1},D_{2},\dots,D_{i}\}$. $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$. We denote the model after the $i$-th incremental training stage as $\theta_{i}=\{\theta_{i}^{agn},\theta_{i}^{spec}\}$, where $\theta_{i}^{agn},\theta_{i}^{spec}$ denote modality-agnostic and modality-specific components, respectively. For MLLM, the modality-agnostic component is the LLM backbone, and modality-specific components include encoders and connectors of each modality.
To mitigate performance degradations in the previously learned modalities, we propose a two-stage MCL paradigm called "MErge then ReAlign" (MERA). In each stage of MCL, MERA executes the following two stages: merging and realigning, to address the forgetting and misalignment respectively. The overall pipeline of MERA is illustrated in Figure 1.
4.1 Stage 1: Merging
Motivation. Model merging has achieved great success in the field of multi-task learning for its superior capability of combining the strengths of multiple models into one. Moreover, Yadav et al. (2024b) finds that model merging is more effective with larger models. Therefore, applying model merging to large-scale models, such as MLLMs, is inherently beneficial. Inspired by these, we introduce model merging to our MCL framework to mitigate forgetting.
However, model merging methods generally can not be directly applied to continual learning. In this work, we only focus on the simplest model merging method, i.e., weight averaging, and revise it into a CL form to provide a basic framework. To adapt weight averaging to CL, we associate its merging coefficients with the progress of CL stages, forming a Cumulative Moving Average (CMA) merging. At the $i$-th training stage, the CMA merged model is calculated by:
$CMA(\theta_{i-1},\theta_{i};i)=\frac{i-1}{i}\theta_{i-1}+\frac{1}{i}\theta_{i}$
To further adapt to the MCL framework, we only merge the modality-agnostic components (the LLM backbone) and ensemble the modality-specific components (encoders and connectors of each modality):
[…]
$\theta_{i-1}^{spec},\theta_{i}^{spec}\}$
To perform MCL, when starting from $\theta_{i-1}$ to incrementally learn a new modality $M_{i}$, firstly we directly train the model on $D_{i}$ to get $\theta_{i,vanilla}$. Note that the training process here is the same as the standard way of training an MLLM on a single modality, except that the initial weights of the model are inherited from $\theta_{i-1}$. Then secondly, we apply CMA merging to $\theta_{i,vanilla}$ and $\theta_{i-1}$ to derive the merged model $\theta_{i,merged}=CMA^{*}(\theta_{i-1},\theta_{i,vanilla};i)$ with integrated knowledge of both new modality and old modalities.
Discussion. Different from other CL methods, the LLM backbone is asynchronously updated with the new connector in the merging stage. To be specific, in the standard MLLM training step, they are synchronously updated indeed, however, in the CMA merging step, only the LLM backbone is updated, resulting in an additional misalignment between them [^2] This is the reason we sample $R_{i}$ from $\{D_{1},D_{2},\dots,D_{i}\}$ rather than $\{D_{1},D_{2},\dots,D_{i-1}\}$ in our realigning stage..
4.2 Stage 2: Realigning
To address the misalignment issue, we simply leverage a small replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$ to further fine-tune all the connectors of $\theta_{i,merged}$ to obtain the final $\theta_{i}$. This process realigns the encoders of each modality with the LLM backbone. The training objective for realigning is unchanged from the original MLLM training objective, i.e., the auto-regressive loss.
@@ 4.2 Stage 2: Realigning
The realigning stage is similar to replay-based CL methods(Scialom et al., 2022; Wang et al., 2024) in form as they both leverage a replay dataset, however, they are essentially different. First, replay methods train on the joint dataset of $D_{i}$ and $R_{i}$, while our realigning stage trains solely on $R_{i}$. Second, the purpose of replay is to review the previous knowledge, while the motivation of realigning stage is to align modality spaces. Third, replay methods update all the parameters to consolidate old knowledge, in contrast, keeping the LLM backbone frozen is crucial in the realigning stage, otherwise the LLM would overfit on replay data.

### v2

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have enhanced their versatility as they integrate a growing number of modalities. Considering the heavy cost of training MLLMs, it is efficient to reuse the existing ones and extend them to more modalities through Modality-incremental Continual Learning (MCL). The exploration of MCL is in its early stages. In this work, we dive into the causes of performance degradation in MCL. We uncover that it suffers not only from forgetting as in traditional continual learning, but also from misalignment between the modality-agnostic and modality-specific components. To this end, we propose an elegantly simple MCL paradigm called "MErge then ReAlign" (MERA) to address both forgetting and misalignment. MERA avoids introducing heavy model budgets or modifying model architectures, hence is easy to deploy and highly reusable in the MLLM community. Extensive experiments demonstrate the impressive performance of MERA, holding an average of 99.84% Backward Relative Gain when extending to four modalities, achieving nearly lossless MCL performance. Our findings underscore the misalignment issue in MCL. More broadly, our work showcases how to adjust different components of MLLMs during continual learning.
@@
Dingkun Zhang1, Shuhan Qi1,2thanks: Corresponding author., Xinyu Xiao1, Kehai Chen1, Xuan Wang1 1Harbin Institute of Technology, Shenzhen 2Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies dingkunzhang0xffff@gmail.com, {shuhanqi,wangxuan}@cs.hitsz.edu.cn, 23b951021@stu.hit.edu.cn, chenkehai@hit.edu.cn
@@ 1 Introduction
With the recent trend of developing general-purpose any-modality Multimodal Large Language Models (MLLMs) (Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024; Shuhan et al., 2023), MLLMs are evolving towards integrating more modalities. The typical MLLM architecture includes modality-specific encoders, modality-specific connectors, and a shared Large Language Model (LLM). A standard process of training MLLMs involves aligning modality-specific components with LLM through modality-text paired data and then fine-tuning on modality-text instruction data (Rao et al., 2024). Such architecture and training strategy have been successfully applied to a wide range of modalities, i.e., image (Liu et al., 2024b, a; Rao et al., 2023), video (Lin et al., 2024a; Maaz et al., 2024), audio (Li et al., 2024; Wu et al., 2024), point cloud (Chen et al., 2024), etc, equipping MLLMs with the ability to understand a growing number of modalities. Existing methods (Wu et al., 2024; Zhan et al., 2024; Panagopoulou et al., 2023; Fu et al., 2024) typically employ a joint training strategy, where the MLLM is jointly trained on datasets of all predefined modalities (Xin et al., 2024). However, it is challenging to extend an existing MLLM to new modalities as it requires another round of joint training on the previous modalities and the new modalities.
To reuse the existing models and adapt them to new data, Continual Learning (CL) is proposed to learn from a stream of data. During continual learning, performance degradation in previously learned tasks often occurs. The degradation is generally attributed to catastrophic forgetting (McCloskey and Cohen, 1989; Goodfellow et al., 2014; Rao et al., 2025), i.e., the model forgets the previously learned knowledge. To this end, many CL methods (Kirkpatrick et al., 2017; Yu et al., 2024b; Scialom et al., 2022; Wang et al., 2024b) have been proposed to alleviate catastrophic forgetting.
In addition to traditional CL, Modality-incremental Continual Learning (MCL) (Yu et al., 2024a) focuses on the particular scenario of incrementally extending MLLMs to new modalities. The exploration of MCL is in its early stages. In this work, we first analyze the causes of performance degradation in MCL. Unlike traditional CL, the performance degradation encountered in MCL comes not only from forgetting but also from the misalignment between modality-agnostic and modality-specific components.
@@
The first stage of MERA aims at addressing the forgetting problem. To this end, we introduce model merging to our MCL framework. We focus on the simplest model merging method, i.e., weight averaging, and revise it into an MCL form. We achieve this by associating its merging coefficients with the progress of CL stages and only merging the modality-agnostic components.
The second stage of MERA aims at addressing the misalignment problem. We leverage a small subset of data from each learned modality to realign the modality encoders with the LLM backbone. In this stage, modality encoders and LLM backbone are both frozen, only the lightweight connectors are updated to enable an efficient realignment between them. Further experiments show that the realigning stage can significantly narrow the gap between the incrementally learned MLLM and the individually trained expert MLLMs on each modality.
@@
  • We analyze the causes of degradation in Modality-incremental Continual Learning (MCL). We uncover that it suffers not only from forgetting as in traditional continual learning, but also from misalignment.
@@
  • Extensive experiments show that our MERA significantly outperforms other representative continual learning methods including the state-of-the-art MCL method, and even achieves nearly lossless MCL performance.
@@
Table 1: Characteristics of different CL categories and our proposed MERA. Extra-train-memory-free: does not introduce extra GPU memory overhead at training time. Arch-modification-free: does not modify the architecture of the model or add auxiliary components. Replay-data-free: does not require access to partial data from the previous tasks or distributions. $\bullet$ denotes that some methods of this category don’t satisfy the property. denotes that this drawback is relatively minor in real applications.
@@ 2.1 Multimodal Large Language Models
Recent advances (Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024) in MLLM have extended LLMs to perceive multimodal inputs such as image, video, audio, point cloud, etc. Among all the MLLMs, the most influential one is LLaVA (Liu et al., 2024b, a), which utilizes a simple MLP connector to project visual information encoded by the pre-trained vision encoder into the language embedding space. Due to its simplicity and effectiveness, LLaVA-like architecture is widely adopted by a wide range of subsequent MLLMs (Lin et al., 2024b, a; Maaz et al., 2024; Wu et al., 2024; Chen et al., 2024) and accounts for the majority of current MLLMs. In this paper, we assume that the MLLM has a LLaVA-like architecture that includes modality-specific encoders and connectors, and a shared modality-agnostic LLM backbone.
The rapid development of MLLMs demands high efficiency in their training process. It is efficient to reuse the existing MLLMs and extend them to more modalities. However, directly fine-tuning MLLMs on new modalities often results in significant performance degradation in previously learned modalities. In this work, we leverage the continual learning technique to tackle this problem.
2.2 Traditional Continual Learning
@@
Continual Learning (CL) (Van de Ven and Tolias, 2019; Wang et al., 2024a) aims to continually acquire new knowledge with minor forgetting of previously learned knowledge. Existing CL methods mainly fall into the following four categories: Regularization-based methods (Kirkpatrick et al., 2017; Huszár, 2017; Schwarz et al., 2018) seek to protect the parameters that store important knowledge. However, storing an importance matrix during training requires extra memory with the same scale as the trainable parameters. Architecture-based methods (Yu et al., 2024a, b; Zadouri et al., 2024; Srinivasan et al., 2023) add task-specific parameters to the base model for each new task. This category requires modifications to the model architecture, harming its reusability. For many methods in this category, the model scale grows linearly as tasks increase, introducing extra memory overhead. Replay-based methods (Rolnick et al., 2019; Scialom et al., 2022; Wang et al., 2024b) leverage a small subset of historical data and replay it when learning on new data. This category requires access to partial data from previous tasks or distributions. However, this drawback is relatively minor in real applications since the replay data is often accessible. Merging-based methods (Wortsman et al., 2022; Marczak et al., 2024; Xiao et al., 2024; Zhu et al., 2024) edit models in parameter space to integrate the previously learned knowledge into the fine-tuned models by model merging (Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a).
Although many traditional CL methods have been proposed, they are not specifically designed for Modality-incremental Continual Learning (MCL). In this work, we propose a simple MCL paradigm tailored for MLLMs. Table 1 summarizes the characteristics of different CL categories and our proposed method.
2.3 Continual Learning for MLLMs
Aside from traditional continual learning, there are methods tailored for MLLMs (Srivastava et al., 2024; Zeng et al., 2024; Maharana et al., 2024; Gao et al., 2024; Cao et al., 2024; Yu et al., 2024a). However, most of these works are specific to vision-language-only MLLMs, and incompatible with MCL where MLLMs can extend to arbitrarily more modalities. To the best of our knowledge, PathWeave (Yu et al., 2024a) is the most relevant work to MCL, and is so far the only work on MCL for MLLMs. It is an architecture-based method that uses an adapter-in-adapter mechanism to alleviate forgetting in previous modalities.
In this work, we dive into the causes of degradation in MCL, uncovering that it suffers not only from forgetting but also from misalignment. To this end, we propose our two-stage MCL paradigm to address both forgetting and misalignment.
Figure 1: Illustration of misalignment and the mechanism of our proposed realigning. $\phi_{i}$ is the feature distribution of the $i$-th modality. Regions in yellow represent the LLM’s expected distribution of the connector’s output. (a) and (b) are the states of the last learning stage and the ideal MCL after learning on a new modality. (c) shows the actual misalignment after learning on a new modality. (d) demonstrates the mechanism of our proposed realigning.
3 Dual Causes of Degradation in MCL
3.1 Preliminary
We define the Modality-incremental Continual Learning (MCL) problem as follows. Given a sequence of $m$ modalities $\{M_{1},M_{2},\dots,M_{m}\}$ and their corresponding datasets $\{D_{1},D_{2},\dots,D_{m}\}$, MCL sequentially learns on each modality $M_{i}$ to obtain the model $\theta_{i}$. We denote the MLLM as $\theta=\{\theta^{Enc},\theta^{Conn},\theta^{LLM}\}$, where $\theta^{Enc},\theta^{Conn},\theta^{LLM}$ denote the modality encoders, the connectors, the LLM backbone, respectively. Notably, the $\theta^{Enc}$ and $\theta^{Conn}$ are modality-specific components and the $\theta^{LLM}$ is a modality-agnostic component. Further, we denote the feature distribution of the outputs of each modality connector $\theta_{i}^{Conn}$ as $\phi_{i}$.
3.2 Forgetting and Misalignment in MCL
@@
MCL is a special scenario of CL, where models incrementally learn on new modalities. However, MCL faces a more severe problem. During continual learning, models would suffer from performance degradation in previously learned domains, tasks, or modalities. In traditional CL, degradation comes from the forgetting of previously learned knowledge. However, in MCL, it comes from two significant aspects: forgetting and misalignment.
Forgetting: the modality-agnostic $\theta^{LLM}$ forgets the knowledge of old modalities. It is associated with various factors, including representation drift (Caccia et al., ), gradient interference (Wang et al., 2023), learning dynamics (Ren and Sutherland, 2024), distribution shift, etc.
Misalignment: the modality-agnostic $\theta^{LLM}$ is misaligned with the modality-specific $\theta^{Enc}$. When incrementally learning on a new modality $M_{i}$, the $\theta^{LLM}$ is updated along with the modality connector of $M_{i}$, while other old connectors are kept frozen. Therefore, $\theta^{LLM}$ adapts to the new feature distribution $\phi_{i}$ and drifts away from the original multimodal distribution $\phi_{1}\cup\phi_{2}\cup\dots\cup\phi_{i-1}$ [^1] For our proposed method, the $\theta^{LLM}$ also drifts away from $\phi_{i}$ due to the use of model merging.. Hence, there exists a misalignment in feature mapping between the old modality encoders and the $\theta^{LLM}$, leading to the breakdown of the "encoder-connector-LLM" collaboration chains for old modalities. The illustration of misalignment is sketched in Figure 1. Experiments in Section 6.1 also provide empirical evidence of the existence of misalignment.
@@
Above, we analyzed the dual causes of performance degradation in MCL. Due to the existence of misalignment, MCL problem requires special treatments in contrast to traditional CL problems.
4 Method
To tackle the dual causes of performance degradation in MCL, we propose a two-stage MCL paradigm called "MErge then ReAlign" (MERA). In each stage of MCL, MERA executes the following two stages: merging and realigning, to address forgetting and misalignment respectively.
4.1 Stage 1: Merging
Model merging is efficient in integrating the previously learned knowledge into the fine-tuned models. Moreover, Yadav et al. (2024b) finds that model merging is more effective with larger models. Therefore, applying model merging to large-scale models such as MLLMs is inherently beneficial. Inspired by these, we introduce model merging to our MCL framework to mitigate forgetting.
The first step before model merging is to perform the standard MLLM training, which often encompasses a pre-training and a fine-tuning phase. After the standard training step, we get the vanilla model $\theta_{i,vanilla}$ that inevitably suffers from forgetting the knowledge of previous modalities.
The second step is to perform model merging to mitigate forgetting. In this work, we only focus on the simplest model merging method, i.e., weight averaging, aiming only to provide a basic framework. To adapt weight averaging to MCL, we associate its merging coefficients with the progress of MCL stages. At the $i$-th training stage, the merged model is calculated by:
$\theta_{i,merged}=\frac{i-1}{i}\theta_{i-1}+\frac{1}{i}\theta_{i,vanilla}$
Notably, we only merge the modality-agnostic component $\theta^{LLM}$ and ensemble the modality-specific components $\theta^{Enc}$ and $\theta^{Conn}$. After merging, we obtain the $\theta_{i,merged}$, whose knowledge of previous modalities is enhanced.
4.2 Stage 2: Realigning
To address the misalignment issue, we propose a lightweight realigning stage to update the multimodal distribution $\phi_{1}\cup\phi_{2}\cup\dots\cup\phi_{i}$ to realign the "encoder-connector-LLM" chains for all the modalities.
The realigning stage simply leverages a small replay dataset [^2] It is different from replay data in replay-based continual learning methods, where their $R_{i}$ is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$. $R_{i}\leftarrow$ sample rr% data from $\{D_{1},D_{2},\dots,D_{i}\}$ to further fine-tune all the connectors of $\theta_{i,merged}$. This realigning process is formulated as:
$\min\limits_{\theta_{i,merged}^{Conn}}\mathbb{E}_{x\sim R_{i}}\mathcal{L}(\theta_{i,merged},x)$
where the $\mathcal{L}$ is the auto-regressive loss, unchanged from the original MLLM training loss. By fine-tuning the lightweight $\theta^{Conn}$ with only a small replay dataset, it efficiently realigns the $\theta^{Enc}$ with $\theta^{LLM}$. After realigning, we can obtain the final model $\theta_{i}$.
@@ 4.2 Stage 2: Realigning
Differences between realigning and replay-based CL methods. The realigning stage resembles replay-based CL methods (Scialom et al., 2022; Wang et al., 2024b) in form as they both leverage replay data, however, they are essentially different. First, replay methods train on the joint dataset of $D_{i}$ and $R_{i}$, while our realigning stage trains solely on $R_{i}$. Second, replay methods update both the $\theta^{LLM}$ and $\theta^{Conn}$, while the realigning stage only efficiently updates the $\theta^{Conn}$ to prevent the knowledge inside $\theta^{LLM}$ from being overwritten again.
4.3 Overall Pipeline
The overall pipeline of MERA is illustrated in Figure 2. In each stage of MCL, MERA goes through two stages. The first stage is merging, where we fine-tune $\theta_{i-1}$ on the incoming modality and merge the fine-tuned model $\theta_{i-1,vanilla}$ with the historical model $\theta_{i-1}$ in order to alleviate forgetting. The second stage is realigning, where we leverage a small set of replay data $R_{i}$ to efficiently fine-tune the lightweight modality connectors $\theta^{Conn}$ to realign the $\theta^{Enc}$ with $\theta^{LLM}$.